Prev Next

SW&IT Meeting Minutes 2022-02-21

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220221
LR20220207 - Bec, Matthieu 80.0
LR20220207 - Mechtley, Brandon 80.0
LR20220207 - Eychaner, Glenn 80.0
LR20220207 - Edmeades, David 80.0
LR20220207 - Grenz, Paul 16.0
LR20220207 - Hooper, Stephen 0.0
LR20220207 - Zhang, Yang 80.0
LR20220207 - Zhang, Xianyu 84.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • archive meeting, LBTI data plans
    • shark-nir AT plans review
    • DMS archive mongo, lit2/webpack
    • ao-confcalib autmoated rev control WIP
    • LN loadrecmat test, issue with rtdb blank fields?
  • This week
    • set_z0 testing (+ explanation from Simone)
    • ao-confcalib autmoated rev control - submit PR and deploy
      • arrange a meeting with AOops folks about some conventions
        • dont use softlink
        • keep scratch unused data in some predefined location like /scratch
    • AO daytime tests setup a wiki location, template, etc.
    • LN loadrecmat follow-up (slack with Alfio)
    • ATT PM

YZ

  • Previous:
    • Worked on a few of the MODS/LBC issues:
      • Ported and updated the "dispMODSx2d" program to Fedora.
      • Reworked certain mods utilities on Fedora.
      • Meeting with Olga on IT#8567 (MODS1R scripts hang), IT#8509 (DOFPIA problem), and some other MODS/LBC issues.
    • Some TMS stuff:
      • Meeting with Heejoo on several software topics.
      • General meeting and some software work.
    • Worked on patching and testing new OSS.
    • Safety training.
  • Next:
    • Continue on some MODS and LBC problems:
      • Working on cleaning old MODS images in /lbt/data/new.
      • Working on cleaning suspicious images on MODS data machines.
      • Continue IT#8567 and IT#8509 discussion with Olga.
    • TMS:
      • Finish remote shutter software update.
      • Possible two tests (remote shutter test, and general software tests).
      • Meeting.
    • Work on releasing the new OSS patch (and setup RPM build for the proxy server).
    • Start email thread with archive team on deploying WCS/new image utilities.

DE

  • Vault seems steady state, probable a few more users will need to upgrade but I'll handle them as they pop up.
  • Need to re-up my Autodesk edu creds.
  • Backups moved to new locations (backup server internal storage and new Synology for Backblaze transfer), going very well. Cleared out some cruft from Backblaze.
    • Need to test Vault restore to verify that relying on Veeam + DB quiescence is valid and we can drop the native Vault DB backup
  • Sanity check backup settings for long horizon DR
  • Finally got SSD that we ordered in November, array rebuilt.
  • RMA SSD is probably in Steward Annex, Cindy will track it down
  • Chasing both Veeam and VMWare licensing, which will be my focus this (short) week.

PSG

  • Last week
    • ARGOS testing, Largly successful
    • Jenkins setup needs some work for proxy.
    • Need to coordinate with Gustavo to switch over ARGOS software from direct connection to use proxy.
    • Continuing work on NOMIC software.
    • Bug investigation of incorrect time causing incorrect filename.
  • This week
    • Test Jenkins Proxy setup
    • Create new GCS branch to implement work there.
    • Finish up NOMIC
    • Test connection to SHARK-NIR computer, try to update landing page.

GE

  • Last week
    • Created a script (startMessages.sh) to open the Messages window on luci.luci.
    • Continued testing the new RTD; see the notes in the pull requests for details, and 8570 (Instrument setup clears the subwindow settings).
    • Worked with SciOps on the LUCI-AO jumps, one cause for which (8578) seems to be PFC being activated after the exposure has started.
    • Continued work on the observer GUI; 6138 (problem with pauses in GUI), 8529 (GUI possibly disconnecting from luci.luci), 8571 (LUCI breezing through offsets).
  • This week
    • Keep testing the new RTD and fix minor bugs as they come up. Fix 8570 (Instrument setup clears the subwindow settings).
    • Report IT 8326 to the Aladin developers.
    • Diagnose and fix the LUCI-AO jumps.
    • Continue work on the observer GUI.
    • Clean up the scripts in lcsp/bin on luci.luci and in stow on observer machines. They have drifted significantly from what is checked into GitHub.
    • Investigate further the black-background (miscolored) LUCI windows; D.Thompson found another instance of the problem using ssh from obs4 to luci.luci.
    • Examine the other LUCI GUIs to check for GUI modifications from threads other than the Event Dispatch Thread.

BM

  • Last week:
    • SOUL-LUCI SW testing, daytime test tutorial from Juan Carlos
      • Logging updates, safeskip button, etc., some debugging of new issues
      • A few more issues:
        • timeouts during WFS boot of AGW devices (have to restart preset)
        • daytime.conf not being loaded automatically,
        • add warning if trying to preset with WFS BCU47 temperature too low (otherwise have to restart preset)
    • Met about SW testing / dev process with GB/MB/XZ
    • Set up second VM for WFS simulation
    • ARGOS testing Friday
  • This week:
    • Catch up on pause/resume/offset problems from the weekend–maybe gain ramp resume is out of order?
    • set_z0 test: schedule with Juan Carlos.
    • ao-confcalib meeting Wednesday (maybe?)
    • Simulate ICE dispatch errors
    • Debug Housekeeper fix
    • ARGOS testing Friday.

SH

  • Installing things on obs for Steve
  • Moving data from old to new Synology
  • VMware installations
    • licenses
  • Telemetry problems

next:

  • VMware PERC cards
  • Moving data from old to new Synology
  • Upgrade 507 computer
Topic revision: r10 - 21 Feb 2022, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback